this composite object is formed by a cone with a hemisphere on top.A formula for the volume of this object is V=1/3(pie) r^3 h+ 2/3(pie) r^3
A) factor this formula
B) use both forms of the formula to calculate the volume of this composite object with radius 4 cm and height 10 cm

Found 3 solutions by ikleyn, stanbon, Alan3354:
Answer by ikleyn(52776)

.
This simple calculation assignment is YOUR job.

Do not ask tutors to make your job.





Answer by stanbon(75887)

this composite object is formed by a cone with a hemisphere on top.
A formula for the volume of this object is V=1/3(pie) r^3 h+ 2/3(pie) r^3
A) factor this formula
V(r) = (1/3)pi*r^3(h+1)
-------------------------------------
B) use both forms of the formula to calculate the volume of this composite object with radius 4 cm and height 10 cm
V(5) = (1/3)*pi^5^3*11)
V(5) = (11/3)*pi*125
-----
V(5) = 458.33*pi cm^3
------------
Cheers,
Stan H.
--------------

Answer by Alan3354(69443)

No math problems with pie, unless it's a pie chart.
---
Use the Greek letter pi.